# Neural Attention Search

### v2

@@ B.1 Collecting the Fine-tune Training Set
To fine-tune  NAtS on LLMs, we collect the training datasets from different tasks:
  • Multi-Document QA: HotPotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and DuReader (zh) (He et al., 2018)
  • Single-Document QA: NarrativeQA (Kociský et al., 2018) and Qasper (Dasigi et al., 2021)
  • Summarization: GovReport (Huang et al., 2021), QMSum (Zhong et al., 2021), MultiNews (Fabbri et al., 2019), and VCSUM (zh) (Wu et al., 2023)
  • Few-shot Leraning: TREC (Li & Roth, 2002), TriviaQA (Joshi et al., 2017), and SAMSum (Gliwa et al., 2019)
  • Code Completion: LCC (Guo et al., 2023) and RepoBench-P (Liu et al., 2024)
Full Duo SLLM H2O NAtS
NarrativeQA 29.2 16.5 21.9 23.6 26.84(14%)
Qasper 41.3 14.6 18.0 18.9 41.05(27%)
MultiFieldQA-en 52.5 29.8 39.9 31.5 53.07(24%)
MultiFieldQA-zh 58.0 29.1 41.6 29.2 55.47(19%)
HotpotQA 49.6 34.9 34.2 42.1 53.02(22%)
2WikiQA 40.0 27.8 31.1 30.4 37.22(25%)
Musique 28.4 13.0 16.4 17.0 26.29(21%)
DuReader (zh) 34.9 26.5 34.4 20.7 34.4(15%)
GovReport 34.9 22.1 30.3 27.1 33.31(20%)
QMSum 25.7 17.2 21.2 20.8 25.48(19%)
MultiNews 27.9 23.4 23.6 23.0 27.32(35%)
VCSUM (zh) 16.3 13.6 14.6 15.1 16.21(13%)
TREC 75.5 52.0 69.0 61.5 73.0(29%)
TriviaQA 88.9 84.8 85.0 86.6 88.8(23%)
SAMSum 47.3 42.1 45.0 45.2 45.68(19%)
LSHT 39.8 17.5 26.0 22.0 39.75(15%)
Passage Count 5.5 3.5 3.5 6.0 4.56(22%)
PassageRetrieval-en 98.0 62.5 30.5 63.0 57.5(19%)
PassageRetrieval-zh 96.5 10.5 22.0 34.5 58.07(17%)
LCC 50.7 48.8 46.4 50.5 51.27(36%)
RepoBench-P 54.8 46.4 49.8 53.8 55.98(30%)
@@
For all the few-shot learning datasets, following  Bai et al. (2024), we randomly concatenate multiple question-answer pairs into one single extended context as one piece of data. The number of concatenated samples for the TREC dataset ranges from$[10,100]$. This value is $[2,6]$ for TriviaQA and $[10,50]$ for SAMSum. Additionally, for the datset that do not have enough context length (e.g., the DuReader dataset), we also merge multiple documents into one piece of data (in our case, this value is 4).
We do not collect all the data whose length goes beyond a threshold to ensure that the context can be fitted into our model. Additionally, we collect at most 500 pieces of data in each dataset since some datasets might not contain pieces of data. In the end, this training set contains 6436 pieces of data in total.
Appendix C Further Results on LongBench Dataset
We show additional results on fine-tuning LLM on the LongBench dataset here. Table 2 shows the results with 25% KV budgets on Mistral-7B-Instruct-v0.3. We found that to achieve the same level of sparsity, Mistral-7B requires a larger sparse regularization value $\lambda$ (the value that we used in Table 2 is $61e-$ 6, which is 10 times larger than the one that we used for Llama3-8B models). The result confirms our conlcusion that  NAtS outperforms the other baselines in most datasets under similar budget level (and many times with an even smaller budgets).
LLAMA3-8B 50% Budget Mistral-7B 50% Budget
Full Duo SLLM H2O NAtS
NarrativeQA 27.3 25.1 24.4 25.26 23.95(20%)
Qasper 28.9 27.8 19.8 21.2 28.74(36%)
MultiFieldQA-en 52.3 50.8 45.6 38.7 50.91(31%)
MultiFieldQA-zh 50.9 52.64 37.3 37.3 48.13(37%)
HotpotQA 40.6 42.45 34.0 38.6 40.76(25%)
2WikiQA 29.3 30.13 25.7 28.1 25.47(27%)
Musique 24.4 22.9 21.6 21.6 23.18(24%)
DuReader (zh) 30.2 29.84 23.2 23.2 26.41(29%)
GovReport 34.1 32.1 32.2 28.5 33.97(29%)
QMSum 24.5 24.61 22.8 22.8 23.63(23%)
MultiNews 27.7 27.8 26.3 25.0 27.92(40%)
VCSUM (zh) 11.3 7.6 13.03 13.0 8.03(28%)
TREC 71.0 73.0 68.0 64.5 69.5(30%)
TriviaQA 87.7 86.9 85.0 85.7 87.36(26%)
SAMSum 42.5 41.37 40.6 40.8 41.32(20%)
LSHT 38.0 31.5 27.2 27.2 37.0(29%)
Passage Count 1.0 1.0 1.0 1.5 5.14(30%)
PassageRetrieval-en 80.5 82.5 42.5 68.5 75.05(22%)
PassageRetrieval-zh 61.2 61.64 51.5 51.5 45.79(32%)
LCC 37.5 37.1 40.3 41.39 38.3(44%)
RepoBench-P 38.0 39.4 41.6 40.0 39.72(35%)
Full Duo SLLM H2O NAtS
NarrativeQA 29.2 26.9 27.3 24.0 28.2(25%)
Qasper 41.3 35.6 25.1 27.9 39.84(40%)
MultiFieldQA-en 52.5 52.5 47.0 40.1 52.8(36%)
MultiFieldQA-zh 58.0 55.4 49.1 39.4 56.26(32%)
HotpotQA 49.6 52.94 42.0 47.2 52.1(34%)
2WikiQA 40.0 39.3 32.8 38.1 39.75(37%)
Musique 28.4 29.45 21.5 20.6 24.3(33%)
DuReader (zh) 34.9 36.44 34.5 26.8 34.42(26%)
GovReport 34.9 32.1 32.6 30.6 33.73(33%)
QMSum 25.7 24.2 22.9 22.5 25.19(31%)
MultiNews 27.9 27.0 26.3 25.6 27.32(47%)
VCSUM (zh) 16.3 15.6 15.7 15.8 16.04(24%)
TREC 75.5 74.5 69.5 66.0 75.0(42%)
TriviaQA 88.9 87.4 88.6 88.4 89.05(35%)
SAMSum 47.3 44.9 46.1 45.8 46.46(29%)
LSHT 39.8 37.5 30.0 31.5 40.5(26%)
Passage Count 5.5 6.0 3.0 4.0 4.5(34%)
PassageRetrieval-en 98.0 99.0 56.0 81.0 94.5(30%)
PassageRetrieval-zh 96.5 96.5 46.5 76.0 87.5(28%)
LCC 50.7 51.1 49.3 52.23 51.84(50%)
RepoBench-P 54.8 53.6 52.3 55.66 54.51(45%)
Table 3: LongBench Results with 50% Budget Size LLama3-8B and Mistral-7B.  .
@@ Appendix C Further Results on LongBench Dataset
Table  3 shows the evaluation results on LongBench with 50% budgets. Despite having fewer KV cache budgets in all the datasets,  NAtS still achieves comparable performance on the LLama3-8B model and better results on the Mistral model and generally performs comparable to the results with the full attention transformers.
Appendix D Ablation Study
D.1 Sparse Regularization Term  $\lambda$
Full NAtS 1e-6 NAtS 5e-7 NAtS 1e-7 NAtS 5e-8 NAtS 1e-8
NarrativeQA 27.3 18.71(3%) 20.79(4%) 23.09(13%) 23.95(20%) 26.11(49%)
Qasper 28.9 25.55(11%) 28.35(13%) 29.01(26%) 28.74(36%) 28.24(63%)
MultiFieldQA-en 52.3 38.76(9%) 45.68(12%) 48.72(22%) 50.91(31%) 53.27(57%)
MultiFieldQA-zh 50.9 40.55(12%) 45.04(15%) 46.03(27%) 48.13(37%) 46.6(64%)
HotpotQA 40.6 31.71(6%) 34.92(8%) 41.7(17%) 40.76(25%) 40.76(49%)
2WikiQA 29.3 25.9(9%) 26.08(11%) 27.39(20%) 25.47(27%) 27.76(51%)
Musique 24.4 16.25(5%) 19.08(7%) 21.0(16%) 23.18(24%) 24.0(48%)
DuReader (zh) 30.2 26.23(7%) 30.0(9%) 28.44(20%) 26.41(29%) 27.69(57%)
GovReport 34.1 25.52(7%) 29.19(9%) 33.47(20%) 33.97(29%) 34.29(55%)
QMSum 24.5 22.88(5%) 23.24(6%) 23.88(15%) 23.63(23%) 24.18(49%)
MultiNews 27.7 26.83(19%) 27.22(21%) 27.77(32%) 27.92(40%) 27.97(65%)
VCSUM (zh) 11.3 12.35(7%) 12.58(9%) 9.4(19%) 8.03(28%) 8.86(58%)
TREC 71.0 57.0(10%) 58.0(11%) 68.5(22%) 69.5(30%) 69.5(56%)
TriviaQA 87.7 84.67(7%) 87.89(9%) 87.42(18%) 87.36(26%) 88.37(51%)
SAMSum 42.5 41.03(6%) 41.39(7%) 41.78(14%) 41.32(20%) 40.91(43%)
LSHT 38.0 23.5(6%) 27.0(8%) 35.5(20%) 37.0(29%) 39.5(58%)
Passage Count 1.0 0.0(8%) 0.5(10%) 1.59(23%) 5.14(30%) 2.09(52%)
PassageRetrieval-en 80.5 12.0(5%) 46.5(7%) 67.0(15%) 75.05(22%) 81.0(47%)
PassageRetrieval-zh 61.2 5.62(10%) 14.88(12%) 41.07(24%) 45.79(32%) 50.34(60%)
LCC 37.5 43.91(18%) 44.07(20%) 40.11(35%) 38.3(44%) 38.5(68%)
RepoBench-P 38.0 42.8(9%) 41.87(11%) 42.63(25%) 39.72(35%) 39.18(59%)

### v3

@@ E.2 Collecting the Fine-tune Training Set
To fine-tune  NAtS on LLMs, we collect the training datasets from different tasks for real-world tasks:
  • Multi-Document QA: HotPotQA [86], 2WikiMultihopQA [35], MuSiQue [76], and DuReader (zh) [34]
  • Single-Document QA: NarrativeQA [46] and Qasper [19]
  • Summarization: GovReport [37], QMSum [94], MultiNews [23], and VCSUM (zh) [80]
  • Few-shot Leraning: TREC [51], TriviaQA [42], and SAMSum [31]
  • Code Completion: LCC [33] and RepoBench-P [54]
We also construct the synthetic passkey-retrieval dataset introduced in DuoAttention [83]. This dataset is generated by embedding multiple random passkeys in different locations with a long context. The model will then be asked to recall this passkey information.
For all the few-shot learning datasets, following  Bai et al. [11], we randomly concatenate multiple question-answer pairs into one single extended context as one piece of data. The number of concatenated samples for the TREC dataset ranges from $[10,100]$. This value is $[2,6]$ for TriviaQA and $[10,50]$ for SAMSum. Additionally, for the datasets that do not have enough context length (e.g., the DuReader dataset), we also merge multiple documents into one piece of data (in our case, this value is 4).
We do not collect all the data whose length goes beyond a threshold to ensure that the context can be fitted into our model. Additionally, we collect at most 500 instances in each dataset since some datasets might not contain data. In the end, the real-world dataset contains 6436 data instances. We add another 564 instances from the synthetic dataset. In the end, our dataset contains 7000 instances. Finetuning a model on this dataset for one epoch takes roughly 8 hours with 2 Nvidia H100 PCIe GPUs.
Appendix F Further Experimental Results
F.1 Results on Ruler Benchmark with Mistral
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
4k 93.68 92.63 58.89 5.60 46.50 57.92 59.33 74.06 44.00 55.73 93.68 (49%)
8k 91.29 89.94 59.21 2.82 38.12 49.30 53.53 83.12 37.66 41.98 90.76 (43%)
16k 89.85 87.51 55.18 3.00 38.59 52.80 63.46 82.83 38.05 34.31 89.09 (41%)
32k 81.24 78.81 47.63 2.70 49.32 66.20 71.91 76.23 46.16 27.92 80.39 (42%)
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
4k 93.68 52.44 42.96 3.26 33.72 38.79 42.74 46.11 33.66 33.26 93.18 (29%)
8k 91.29 40.96 36.54 2.37 31.36 36.67 42.01 46.73 28.64 25.47 89.97 (24%)
16k 89.85 35.42 33.38 1.92 30.30 37.48 42.84 57.17 28.17 25.55 87.69 (21%)
32k 81.24 40.00 32.76 1.70 35.17 44.46 60.56 61.55 31.84 24.88 78.60 (21%)
Table 3: Ruler results with 50% KV budgets (top) and 25% KV budget (bottom) on Mistral-7B-Instruct-v0.3. All Full models use the 100% KV budgets. We mark the actual KV budgets used by  NAtS in the bracket.
Table  3 shows the results of different optimizers on the Mistral-7B-Instruct-v0.3 model. In this case, the $\lambda$ for 25% and 50% budgets are 1e61e-6 and 3e73e-7 respectively. We found that to achieve the same sparsity level, Mistral-7B requires a larger sparse regularization value $\lambda$ compared to the Llama 3.1 model. This might indicate that the attention map values in Mistral are more evenly distributed, and thus we need a larger $\lambda$ to force more tokens to only focus on the local information. The maximal context length of Mistral 7B is $32k$. We only evaluate the models with a maximal length of $32k$. The result is consistent with the results shown in Table 1.  NAtS consistently outperforms the other baselines with smaller budgets.
F.2 Further results on LongBench
We show additional results on fine-tuning LLM on the LongBench dataset here. Table 4 shows the results with 25% KV budgets on Mistral-7B-Instruct-v0.3 with $\lambda=3e-6$. The result confirms our conclusion that  NAtS outperforms the other baselines in most datasets under a similar budget level (and many times with even smaller budgets).
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
NarrativeQA 29.21 16.49 27.56 23.65 22.70 24.34 21.22 27.72 21.97 21.67 26.7(9%)
Qasper 41.34 14.58 29.16 18.95 23.08 22.99 21.21 28.85 23.54 19.72 41.46(21%)
MultiFieldQA-en 52.51 29.83 31.88 31.52 38.56 38.11 34.11 45.18 35.34 28.80 52.36(18%)
MultiFieldQA-zh 58.03 29.15 30.47 29.21 31.42 32.30 32.40 38.12 29.67 23.20 54.13(16%)
HotpotQA 49.62 34.85 46.07 42.10 46.37 48.83 46.93 47.48 45.94 39.64 51.45(15%)
2WikiQA 40.01 27.83 35.11 30.43 33.22 33.14 32.03 38.24 33.96 27.02 37.05(18%)
Musique 28.44 13.03 22.65 16.99 22.45 21.28 22.62 25.01 23.67 16.22 27.88(14%)
DuReader (zh) 34.87 26.51 18.69 20.66 25.34 27.55 28.45 27.04 24.82 19.11 35.25(12%)
GovReport 34.94 22.09 27.44 27.11 28.31 27.94 28.94 29.33 26.94 25.58 33.39(14%)
QMSum 25.67 17.25 22.11 20.84 21.43 21.95 21.95 22.77 22.39 19.53 24.70(13%)
MultiNews 27.85 23.41 22.01 23.03 23.77 23.62 23.34 24.37 23.89 22.55 27.35(28%)
VCSUM (zh) 16.34 13.58 15.41 15.13 15.28 15.38 14.69 16.10 14.96 14.87 16.06(11%)
TREC 75.50 52.00 70.50 61.50 58.50 59.00 57.00 67.50 58.00 63.00 73.50(20%)
TriviaQA 88.89 84.85 88.92 86.60 88.89 89.39 89.19 88.98 89.31 86.94 89.80(17%)
SAMSum 47.29 42.15 45.44 45.16 46.40 46.72 46.51 46.67 46.58 43.90 46.91(13%)
LSHT 39.75 17.50 29.00 22.00 38.50 39.00 39.50 39.25 37.50 17.25 38.5(11%)
Passage Count 5.50 3.50 6.00 6.00 6.00 8.00 5.50 6.00 6.50 5.00 4.5(14%)
PassageRetrieval-en 98.00 62.50 86.50 63.00 91.50 94.00 87.50 89.50 88.00 17.25 96.00(13%)
PassageRetrieval-zh 96.50 10.50 26.50 34.50 69.25 83.25 71.00 90.00 75.50 16.00 92.00(14%)
LCC 53.02 50.87 51.22 52.77 55.67 55.67 53.76 56.17 55.46 51.08 53.81(29%)
RepoBench-P 56.83 48.93 54.80 55.74 56.82 57.06 56.75 56.75 56.40 53.59 56.7(23%)
@@ F.2 Further results on LongBench
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
NarrativeQA 31.35 29.40 28.18 20.15 30.37 30.04 29.49 32.03 29.18 13.92 31.39(29%)
Qasper 24.73 20.37 18.15 18.09 18.44 19.42 17.24 22.13 18.54 22.25 25.17(45%)
MultiFieldQA-en 29.46 26.70 17.89 22.50 23.21 24.62 22.55 26.30 22.95 14.24 28.94(41%)
MultiFieldQA-zh 60.01 60.31 42.19 38.88 47.66 50.23 44.54 56.22 46.50 23.91 61.27(37%)
HotpotQA 17.06 18.84 16.73 15.61 16.55 16.29 17.22 15.83 15.86 8.85 17.28(37%)
2WikiQA 16.64 16.30 14.28 11.71 15.82 15.38 14.67 17.01 14.81 10.58 16.84(41%)
Musique 11.59 13.77 10.75 10.12 10.78 11.19 10.63 12.85 10.64 5.03 11.89(36%)
DuReader (zh) 35.56 33.29 18.16 27.32 29.22 30.82 28.49 32.81 27.93 22.80 33.36(32%)
GovReport 34.30 32.99 30.84 29.46 31.34 31.51 31.92 33.09 30.54 25.45 34.23(36%)
QMSum 23.30 23.89 21.43 20.72 22.08 22.74 22.22 22.89 22.92 19.76 22.6(33%)
MultiNews 27.13 26.29 24.77 25.10 25.27 25.48 25.08 25.81 25.10 24.86 27.17(51%)
VCSUM (zh) 16.36 15.70 15.29 15.03 15.70 15.77 15.59 15.89 15.32 14.67 16.24(28%)
TREC 72.50 72.50 71.00 65.50 60.00 66.50 63.50 70.50 59.50 64.00 73.50(44%)
TriviaQA 91.15 90.41 91.44 90.25 91.47 90.97 91.63 91.14 90.98 81.17 92.14(40%)
SAMSum 43.72 42.68 43.41 42.57 44.21 43.44 43.00 43.41 43.91 39.64 42.61(30%)
LSHT 46.50 46.50 40.00 34.50 45.00 46.00 45.00 46.50 45.00 24.00 45.5(31%)
Passage Count 6.63 6.67 6.73 2.21 6.33 6.07 6.03 4.74 8.04 1.80 9.78(37%)
PassageRetrieval-en 97.98 98.55 96.67 93.54 95.98 97.88 95.83 98.12 96.10 17.96 97.32(36%)
PassageRetrieval-zh 77.99 75.58 43.61 67.83 79.66 78.27 72.38 78.22 79.48 19.72 78.17(34%)
LCC 54.10 56.22 53.10 55.61 54.08 53.13 54.23 52.33 51.87 52.28 53.27(57%)
RepoBench-P 51.39 57.54 50.80 55.17 52.49 51.59 52.32 52.38 51.59 48.46 52.68(47%)
Table 5: LongBench Results with 50% Budget Size on LLama 3.1 8B .
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
NarrativeQA 29.21 26.95 28.76 24.00 24.65 25.40 25.66 28.20 24.41 23.58 28.1(31%)
Qasper 41.34 35.65 35.80 27.88 31.68 31.94 31.38 38.10 31.50 28.60 41.75(46%)
MultiFieldQA-en 52.51 52.50 37.35 40.15 46.12 48.52 44.73 52.20 45.70 35.52 52.79(42%)
MultiFieldQA-zh 58.03 55.37 36.43 39.43 40.55 40.92 43.70 47.93 38.24 27.48 56.75(38%)
HotpotQA 49.62 52.94 47.66 47.24 49.34 49.57 49.51 47.20 51.26 39.64 50.73(39%)
2WikiQA 40.01 39.26 38.44 38.09 37.27 36.75 40.24 38.13 37.49 31.42 40.77(42%)
Musique 28.44 29.45 27.24 20.61 26.43 25.77 27.28 28.54 25.85 16.84 28.64(38%)
DuReader (zh) 34.87 36.44 18.63 26.82 30.29 32.18 32.72 32.18 29.71 20.79 34.7(33%)
GovReport 34.94 32.07 30.79 30.58 31.43 31.55 32.16 32.39 29.63 25.80 34.23(38%)
QMSum 25.67 24.20 22.91 22.48 23.16 24.36 23.47 25.12 24.11 21.08 25.94(37%)
MultiNews 27.85 26.97 25.18 25.62 25.97 26.24 25.83 26.26 25.56 24.39 27.47(54%)
VCSUM (zh) 16.34 15.58 16.24 15.80 15.94 16.42 15.73 16.57 15.63 14.13 16.55(31%)
TREC 75.50 74.50 74.00 66.00 66.50 68.50 69.00 74.50 64.50 69.00 75.00(49%)
TriviaQA 88.89 87.37 89.06 88.36 88.71 88.81 88.69 88.08 89.55 87.05 89.3(41%)
SAMSum 47.29 44.85 47.30 45.82 47.19 47.30 46.77 47.31 46.85 45.29 46.87(35%)
LSHT 39.75 37.50 33.00 31.50 39.25 39.25 39.25 40.25 38.00 21.00 38.75(32%)
Passage Count 5.50 6.00 5.50 4.00 4.00 5.50 5.00 5.00 4.00 4.50 7.50(39%)
PassageRetrieval-en 98.00 99.00 89.50 81.00 97.50 98.50 97.00 98.00 98.50 33.00 98.0(35%)
PassageRetrieval-zh 96.50 96.50 51.50 76.00 89.75 95.50 92.00 97.00 96.50 17.50 96.0(35%)
LCC 53.02 53.22 52.94 54.40 54.50 54.14 53.79 54.63 53.78 52.04 53.66(56%)
RepoBench-P 56.83 55.89 55.47 57.55 56.76 57.05 55.68 56.38 56.36 54.67 57.38(52%)
Table 6: LongBench Results with 50% Budget Size on Mistral-7B. .
@@
Tables  5 and  6 show the evaluation results on LongBench with 50% budgets. Despite having fewer KV cache budgets in all the datasets,  NAtS still achieves comparable performance on the LLama3-8B model and better results on the Mistral model and generally performs comparable to the results with the full attention transformers.
Appendix G Ablation Study
G.1 Sparse Regularization Term  $\lambda$
Full NAtS 1e-6 NAtS 5e-7 NAtS 3e-7 NAtS 1e-7 NAtS 5e-8
NarrativeQA 31.4 27.05(5%) 29.19(9%) 30.53(13%) 31.39(29%) 31.94(41%)
Qasper 24.7 31.39(15%) 28.6(21%) 23.76(27%) 25.17(45%) 26.97(58%)
MultiFieldQA-en 29.5 26.84(13%) 27.41(18%) 28.94(24%) 28.94(41%) 29.7(54%)
MultiFieldQA-zh 60.0 51.73(14%) 56.64(17%) 61.18(22%) 61.27(37%) 61.87(49%)
HotpotQA 17.1 15.47(10%) 17.44(15%) 16.06(20%) 17.28(37%) 16.65(49%)
2WikiQA 16.6 15.37(13%) 16.44(18%) 17.05(23%) 16.84(41%) 16.75(53%)
Musique 11.6 10.78(9%) 9.87(13%) 11.42(18%) 11.89(36%) 11.31(48%)
DuReader (zh) 35.6 30.06(9%) 33.63(13%) 34.28(17%) 33.36(32%) 34.41(45%)
GovReport 34.3 30.82(10%) 32.47(14%) 33.82(19%) 34.23(36%) 34.8(50%)
QMSum 23.3 23.37(8%) 23.09(12%) 23.11(17%) 22.6(33%) 22.74(45%)
MultiNews 27.1 26.45(22%) 26.57(28%) 26.72(33%) 27.17(51%) 27.11(64%)
VCSUM (zh) 16.4 15.56(7%) 15.61(10%) 15.61(14%) 16.24(28%) 15.99(39%)
TREC 72.5 68.0(14%) 70.5(20%) 72.0(26%) 73.5(44%) 73.0(56%)
TriviaQA 91.2 91.38(12%) 90.32(17%) 91.61(22%) 92.14(40%) 91.64(52%)
SAMSum 43.7 43.98(8%) 44.0(12%) 44.21(16%) 42.61(30%) 43.84(41%)
LSHT 46.5 46.0(8%) 44.0(12%) 47.5(16%) 45.5(31%) 45.5(43%)
Passage Count 6.6 2.56(10%) 6.88(14%) 7.12(20%) 9.78(37%) 8.82(50%)
PassageRetrieval-en 98.0 48.17(9%) 82.65(14%) 95.81(19%) 97.32(36%) 96.93(48%)
PassageRetrieval-zh 78.0 33.14(12%) 64.77(16%) 78.5(20%) 78.17(34%) 77.7(46%)
LCC 54.1 55.46(24%) 54.53(31%) 53.54(38%) 53.27(57%) 54.4(69%)
RepoBench-P 51.4 50.78(14%) 51.85(21%) 53.48(28%) 52.68(47%) 52.36(59%)